INAF queue observing, 2017 Apr 14-15 
Observer: R. Carini, M. Faccini 
Telescope Operator:  S. Allanson 
Support Astronomer: O. Kuhn  
Instrument:  MODS; LUCI
  Summary 
At the sunset clear
Night duration (12\xB0 el twilight)  03:00am-11:50pm UT) = 
08 h 50m
Observing time: 
08 hrs 20 m (94%)
Weather time loss: 
0 hrs  (0%)
technical time loss 
30 m (%) the disconnection was about 3h, but the TO has observed for us form the mountain,  during this period LUCI was stucked a lot of time, the time lost is not evaluable.
  Data Summary 
	
		
			|  Project_RUN  | 
			 Objects  | 
			 Filter  | 
			 ExpTime  | 
			 Notes  | 
		
	
	
		
			|  SPecPhot star  | 
			 Feige34  | 
			 blue grating MODS1; red grating MODS2, flat g and r in MODS1+MODS2  | 
			    | 
			    | 
		
		
			|  Palazzi  | 
			 ATLAS  | 
			 blue grating MODS1 ; red grating MODS2  | 
			 20m img + 2h spec  | 
			 completed  | 
		
		
			|  Ricci A2  | 
			 NGC2712  | 
			 J  | 
			 300s LUCI1+ 300s LUCI2  | 
			 airmass out of spec, no telluric  | 
		
		
			|  Prandoni  | 
			 USS439  | 
			 K  | 
			 20m LUCI1+20m LUCI2  | 
			 completed  | 
		
		
			|  Prandoni  | 
			 USS32  | 
			 K  | 
			 20m LUCI1 + 20m LUCI2  | 
			 completed  | 
		
		
			|  Prandoni  | 
			 USS320  | 
			 K  | 
			 20m LUCI1 + 20m LUCI2  | 
			 completed  | 
		
		
			|  Moscadelli J  | 
			 g075  | 
			 H2  | 
			 1h 30m  | 
			 completed  | 
		
		
			|  Standard  | 
			 F075  | 
			 H2  | 
			    | 
			    | 
		
		
			|  Calibration LUCI  | 
			 sky flats  | 
			 H2, K  | 
			    | 
			    | 
		
		
			|  Calibration MODS  | 
			    | 
			    | 
			    | 
			    | 
		
	
  Details 
00:31 test images for MODS1 and MODS2 
-  
mods1r.20170414.0072, mods1b.20170414.00--
  -  
mods2r.20170414.0068, mods2b.20170414.00--
 
 
00:40 calibration with MODS for Palazzi program
02:27 Acq standard Feige34 for MODS1 and MODS2 
-  
mods1r.20170414.0073
  -  
mods2r.20170414.0069
 
 
02:37 ModsAlign  standard Feige34 for MODS1 and MODS2 
-  
mods1r.20170414.0074
  -  
mods2r.20170414.0070
 
 
02:47 Science  standard Feige34 for MODS1 and MODS2 (seeing 0,9) 
-  
mods1b.20170415.0001 - 0003
  -  
mods2r.20170415.0001 - 0003
  -  some communication glitch with MODS2, we manually stop the first exposure
 
 
03:05  switch to Palazzi program, target Atlas 
-  execBinomods atlas_img.img
  -  
mods1b.0004 - 0006; mods1r.0001-0003
  -  
mods2b.0001-0003; mods2r.0004-0006     seeing 0,9 the objects is near the intersection of the quadrants but it's easy to see it
 
 
03:27  switch to spectroscopy program, target Atlas 
-  acqBinoMods atlas.acq
  -  
mods1r.0004-0006
  -  
mods2r.0007-0009     seeing 0,9
 
 
03:41  modsAalign
03:59  the object is too faint to be seen, we manually change the exposure time to 5 min to check again 
 
04:06  modsAlign -y 7
04:13  execBinomods atlas 
-  
mods1b.0007-0010
  -  
mods2r.00011-0014
 
 
04:31 seeing variable. the mean seeing is about 1, but there are some peaks to 1.3'' or 1.7'' 
-  05:24 seeing 0.8''-0.9''
  -  06:00 SEEING 1.2"
  -  06:06 SEEING 0.9''
 
 
06:40 execBinomods  atlas 
-  
mods1b.0007-0010
  -  
mods2r.00011-0014
 
 
06:30 switch to LUCI
06:45 Ricci program, target A2. GO FRicci_1617_20_A2_newguide_bino.xml 
-  acquisition 
luci1.20170415.0006-0009 , luci2.20170415.0005-0008
  -  science 
luci1.0010- , luci2.0009-001
  -  airmass out of spec
  -  07:19 lost connection with mountain, luci is running but do not see anything
  -  07:28 luci stucked
 
 
07:48 the TO try to run the script from the mountain. We decide to slew to Prandoni program because is an image program and the TO could easier run the script. 
-  The telescope is configured for the bino mode, so we use the bino script but with the same exposure time, to facilitate the TO and to be faster.
  -  load and GO USS439.xml in twin version. Moreover
  -  from here we can not see anything, I do not know the number of the image, how much is the seeing, the quality and so on.
  -  the TO said that the seeing is about 0.7''-0.8''
  -  luci has problem, stuck many times
  -  08:45* the script stopped. the TO run again the script
  -  08:49* luci stuck again. the TO run again the script
 
 
08:53 load and GO  USS32.xml 
-  luci stuck again
  -  seeing 0.7''-0.8''
 
 
09:36 load and GO USS320.xml 
-  10:14 script stuck. the TO run again the script
 
 
10:23 the problem of the connection is resolved!!!!!!!! EUREKA!!!!!
10:26 slew to Moscadelli program, target B 
-  GO g075_H2_bino.xml
  -  problem with the preset but finally it works!!!
  -  
luci1.0089-0106 , luci2.0087-0104
  -  seeing about 1.0''
  -  11:23 seeing 0.7''-0.8''
  -  11:37 seeing about 0.6''
 
 
11:47 we stop the acquisition to slew to the Standard F149 
 
11:48 slewing to Standard for g075
   *last 2 images elongated
12:02 filter K - elongated also in K
12:14 we collimate on a nearby star and take the filters again
  *there are other troubles like "jumps features" 
luci2.0087-0117
12:31  we slew to a blank field to make the calibration in K
   
luci1.0122-0127 good, 
luci2.00122-0127 good
12:44 second script in H2
  
luci1.0128-130 , 
luci2.00134-0138
12:50 closing the dome to make MODS calibration 
-  flat slitless execBinoMods grpixflats_b.cal grpixflats_r_mods2.cal 
mods1b.0011-0020 , mods2r.0015-0019
  -  flat with slit 1.2'' 
mods1b.0021-0026 , mods2r.0020-0025
  -  flat with slit 5.0'' 
mods1b.0027-0032 , mods2r.0026-0028
  -  we test some flats without slit for MODS1 blue channel 
mods1b.0033-0036 are not good
  -  
mods1b.0037-0041 VFlat 4 lamp with texp =3 are good
  -  other test 
mods1b.0042-0043
  -  arc 
mods1b.0044-0046 , mods2r.0029-0031
  -  flat for imaging 
mods1r.0008-0012, mods1b.0047-0051 , mods2r.0032-0036, mods2b.0004-0008
  -  bias dual 3k 
mods1r.0013-0017, mods1b.0052-0056 , mods2r.0037-0041, mods2b.0009-0013
  -  bias dual 8k 
mods1r.0018-0022, mods1b.0057-0061 , mods2r.0042-0046, mods2b.0014-0018
 
 
internal flat fields, gr and ui
-- %{RobertaCarini - 2017-04-14}%